A HIGH RESOLUTION STUDY OF THE ELECTRONIC SPECTRA OF CONFORMERS OF 1,3 AND 1,4 DIVINYLBENZENE: UNEXPECTED EXCITED STATE BEHAVIOR^{a}

\underline{TRI NGUYEN}, JASON W. RIBBLETT, and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260.

The vibrationally resolved S_{1}←S_{0} electronic spectra of 1,3- and 1,4-divinylbenzene each exhibited two origin bands that are ˜ 800 cm^{−1} and ˜ 1000 cm ^{−1} apart respectively. Intrigued by this observation, we have recorded the different bands at full rotational resolution in a molecular beam. Analyses of these data show that the two bands are the S_{1}←S_{0} of the different conformers, cis and trans, in each molecule, which are remarkably far apart. Typically, such systems exhibit more closely spaced origin bands. A model that accounts for this unexpected behavior will be discussed.

awork supported by NSF.